OpenStudy (anonymous):

Helena made trail mix and wants to list the ingredients in order of increasing weight. She used one and five eighths lb of mixed nuts, five eighths lb of chocolate covered raisins, and one and two eighths lb of dried mixed fruit. If she decides to include one and one half lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight?

OpenStudy (anonymous):

so there's \(\large 1+\frac{ 5 }{ 8 } ~and~ 1+\frac{ 2 }{ 8 } ~and~\frac{ 5 }{ 8 }~and~1+\frac{ 1 }{ 2 }\)

OpenStudy (anonymous):

everything is pretty much written as a fraction of 8 except the sunflower seeds. So it would probably be the easiest to just rewrite that as a fraction of 8 and than it would be easy to sort them.

OpenStudy (anonymous):

a. seeds, nuts, raisins, fruit b. nuts, seeds, raisins, fruit c. raisins, fruit, nuts, seeds d. raisins, fruit, seeds, nuts

OpenStudy (anonymous):

I am not very sure I:

OpenStudy (anonymous):

if you write 1/2 as a fraction of 8 what would it be? \[\frac{ 1 }{ 2 }=\frac{ ? }{ 8 }\]

OpenStudy (anonymous):

4/8?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

so now we have \[ \ 1+\frac{ 5 }{ 8 } mixed ~nuts \] \[ 1+\frac{ 2 }{ 8 } ~mixed~fruit\] \[\frac{ 5 }{ 8 } raisins\]\[1+\frac{ 4 }{ 8 } sunflower \]

OpenStudy (anonymous):

So d? O: :)

OpenStudy (anonymous):

all you have to do i see which one is larger than the other. and sort them from biggest to smallest

OpenStudy (anonymous):

Thank you !!!!!!!! :)

OpenStudy (anonymous):

yw